On January 1, 2020, Teal Mountain Co. leased a building to Sandhill Inc. The relevant information related to the lease is as follows. Prepare the journal entries that Teal Mountain should make in 2020. Prepare the journal entries that Sandhill should make in 2020. If Sandhill paid $40,000 to a real estate broker on January 1, 2020, as a fee for finding the lessor, what is the initial measurement of the right-of-use asset?

1. The lease arrangement is for 10 years. The building is expected to have a residual value at the end of the lease of $3,700,000 (unguaranteed).
2. The leased building has a cost of $4,200,000 and was purchased for cash on January 1, 2020.
3. The building is depreciated on a straight-line basis. Its estimated economic life is 50 years with no salvage value.
4. Lease payments are $275,000 per year and are made at the beginning of the year.
5. Sandhill has an incremental borrowing rate of 6%, and the rate implicit in the lease is unknown to Sandhill.
6.

Both the lessor and the lessee are on a calendar-year basis.
Answer #1

Part A

Date

Account titles and explanation

Debit

Credit

1/1/20

Buildings

470000

Cash

470000

12/31/20

Cash

275000

Rent revenue

275000

Depreciation Expense

94000

Accumulated Depreciation—Buildings (470000/50)

94000

Part B

Date

Account titles and explanation

Debit

Credit

12/31/20

Rent expense

275000

Cash

275000

Part C

Initial measurement of the right-of-use asset = 2145465+4000 = $2149465

Present value = 275000*present value of annuity due of $1 @ n= 10, i=6% = 275000*7.80169 = $2145465

40000/10 = 4000